# Neural ODEs for holographic transport models without translation symmetry

### v2

@@ Appendix C At low temperatures
We will change the previous ansatz into
@@
and
@@
The former inputs the temperature $T$, which reduces the difficulty of learning near-horizon geometry. The latter is beneficial for learning the mass square in the GA model at low temperatures, where $m^{2}/z^{2}$ grows fast near the horizon. Note that for the low-temperature MG and LA models, as well as all the relatively high-temperature models studied in the main text, the ansatz (A.28) remains applicable, and the performance of Neural ODEs will not be significantly affected. The only price of using eq. (A.28) is to assume that the mass square is positive [^11] Although there is no general argument, the mass square in all the models given in Hartnoll1601 is positive, and the absence of instabilities actually requires the positive mass square in MG and GA models Vegh1301 ; Baggioli2014 ..
@@
When learning the mass square of MG and LA models, we initialize the neural network with $\mathcal{N}(0,0.001)$. For other learning targets, we use $\mathcal{N}(0,1)$.
In figures 7-10 and table 2, we generate the data and illustrate the performance of Neural ODEs at low temperatures. We choose the parameters ($5.92\alpha=$ 5.9, $\beta^{2}=5.9$, $\gamma_{1}=4.9$, $\gamma_{2}=1$), which correspond to the same temperature $T\simeq 0.004$.
Figure 7: Training data and performance of Neural ODEs for the MG model at low temperatures.
Figure 8: Training data and performance of Neural ODEs for the LA model at low temperatures.
Figure 9: Training data and performance of Neural ODEs for the GA model at low temperatures.
Figure 10: The RMSE for evaluating the hyperparameter $b$ in three models at low temperatures.
@@ Appendix D Entanglement entropy
With these preparations in place, let’s substitute the bulk metric (the learned metric or the true value) into eq. (A.31) and get $l(z_{\ast})$ by numerical integration. By the interpolation, we can construct the inverse function $z_{\ast}(l)$. In terms of eq. (A.34) with $z_{\ast}(l)$, the derivative $S^{\prime}(l)$ can be obtained. In figure (11) and figure (12), we plot $S^{\prime}(l)$ for the three models at (relatively) high temperatures and low temperatures. We also calculate the MRE by uniformly sampling 10000 points within the allowed range of $l$, see table (3).
Figure 11: The performance of Neural ODEs in predicting $S^{\prime}(l)$. From left to right, the three panels represent the results for the MG, LA and GA models at high temperatures.
Figure 12: The performance of Neural ODEs in predicting $S^{\prime}(l)$ for three models at low temperatures.

### v3

@@ Appendix C At low temperatures
Figure 7: Training data and performance of Neural ODEs for the MG (top), LA (middle) and GA (bottom) model at low temperatures.
Figure 8: The hyperparameter turning of $2\pi T$ (top) and $b$ (bottom) in three models at low temperatures.
@@
We change the previous metric ansatz as
@@
where the temperature $T$ will be specified by hyperparameter tuning. Note that this scheme reduces the difficulty of learning near-horizon geometry.
We also change the previous mass ansatz as
@@
This is beneficial for learning the mass square in the GA model at low temperatures, where $m^{2}/z^{2}$ grows fast near the horizon. Note that for the low-temperature MG and LA models, as well as all the relatively high-temperature models studied in the main text, the ansatz (A.29) remains applicable, and the performance of Neural ODEs will not be significantly affected. The only price of using (A.29) is to assume that the mass square is positive [^14] Although there is no general argument, the mass square in all the models given in Hartnoll1601 is positive, and the absence of instabilities actually requires the positive mass square in MG and GA models Vegh1301 ; Baggioli2014 ..
@@
When learning the mass square of MG and LA models and tuning the hyperparameters $T$ and $b$, we initialize the neural network with the normal distribution $\mathcal{N}(0,0.001)$. For other tasks, we use $\mathcal{N}(0,1)$.
6) Hyperparameter tuning
The method of tuning $b$ remains the same as before. For tuning $T$, the main difference is in the dataset separation: we randomly shuffle the dataset within the frequency range $[0.01,\omega_{\max}]$ and split it into two parts: 70% for the training set and 30% for the validation set. Additionally, we remove the L1 regularization.
In figures 7-8 and table 2, we generate the data and illustrate the performance of Neural ODEs at low temperatures. We choose the parameters ($5.92\alpha=$ 5.9, $\beta^{2}=5.9$, $\gamma_{1}=4.9$, $\gamma_{5}=1$), which correspond to the same temperature $T\simeq 0.004$.
@@ Appendix D Entanglement entropy
Figure 9: The performance of Neural ODEs in predicting $S^{\prime}(l)$. From left to right, the three panels represent the results for the MG (left), LA (middle) and GA (right) models at high (top) and low (bottom) temperatures.
Figure 10: Relative errors in $S^{\prime}(l)$ for three models at high (top) and low (bottom) temperatures.
@@
With these preparations in place, let’s substitute the bulk metric (the learned metric or the true value) into eq. (A.32) and get $l(z_{\ast})$ by numerical integration. By the interpolation, we can construct the inverse function $z_{\ast}(l)$. In terms of eq. (A.35) with $z_{\ast}(l)$, the derivative $S^{\prime}(l)$ can be obtained. In figure (9), we plot $S^{\prime}(l)$ for the three models at (relatively) high temperatures and low temperatures. The change of the relative error along $l$ is depicted in figure (10). Notably, the relative error decreases oscillatingly as $l$ increases. At last, we calculate the MRE by uniformly sampling 10000 points within the allowed range of $l$, which are given in Table 3.